babyboi23:

In Ogilvie's class of 40 students, 24 students report that English is their favorite class and 40% of the students report that keyboarding is their favorite class. Show your work to explain how to determine whether more students prefer English or keyboarding.

HelperA:

40/24 is 60 percent, if you were to divided 40/4=10*6=60% For English, so in that case English should win by 60%/40%

babyboi23:

thanks

HelperA:

Can I have a medal?

babyboi23:

yhea

HelperA:

You rock, Let me know If you need anything else.

babyboi23:

cool man